OpenStudy (anonymous):

Solve each equation, showing all steps leading to your answer. Then check your answer to determine if any of the solutions are extraneous 1. x+1−−−−−√+5=x 2. 5x−−√−x=0 3. 2x+1−−−−−√+1=x 4. 3x−2−−−−−√+3=4x

OpenStudy (anonymous):

\[\sqrt{x+1}+5=x\] \[\sqrt{5x}-x=0\] \[\sqrt{2x+1}+1=x\] \[\sqrt{3x-2+3=4x}\] Here are the equations

OpenStudy (anonymous):

i'll do the first one others are similar \[\sqrt{x+1}+5=x\] subtract \(5\) \[\sqrt{x+1}=x-5\] square both sides (correctly) \[x+1=(x-5)^2\] \[x+1=x^2-10x+25\] solve the quadratic equation by setting it equal to zero and factoring \[x^2-11x+24=0\] \[(x-3)(x-8)=0\] set the factors equal to zero and get \[x-3=0\iff x=3\] or \[x-8=0\iff x=8\] then check the answer in the original equation

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

\[\sqrt{3+1}+5=3\] \[\sqrt{4}+5=3\] \[2+5=3\] no not true \(3\) is not a solution \[\sqrt{8+1}+5=8\] \[\sqrt{9}+5=8\] \[3+5=8\] true, \(8\) is a solution

OpenStudy (anonymous):

Thanks! I appreciate it a lot!!

OpenStudy (anonymous):

yw